write an equation of the line that is parallel to the line y=x+3 and passes through the point (1,2)


use thes ordered pairs to find the slopes of the line AB and line CD. Also tell whether the lines are parallel, perpendicular or neither.
A(-1,-4), B(2,11), C(1,1),D(4,10)

Answer by ewatrrr(24785)

 
Hi,
line containing the given point(1,2) and parallel to y=x+3
y = x + b ||Using Pt(1,2) to solve for b
2 = 1 + b
1 = b
y = x + 1
A(-1,-4),
B(2,11) m = -15/-3 = 5+
C(1,1),
D(4,10) m = -9/-3 = 3
Neither